A water heater is designed to last an average of 16 years,the lifetime of such a heater can be modelled by an exponential random variables X with lambda value of 0.0625, suppose a heater is selected at random ,,,

What is the probability the heater will last at least 5 years


Suppose the heater last for 5 years what is the probability it will last for the other 5 years


1
Expert's answer
2021-12-08T11:15:14-0500

Let "X=" the lifetime of a heater: "X\sim Exp(\lambda)."

Given "\lambda=1/16=0.0625"

i.


"P(X\geq 5)=e^{-\lambda(5)}=e^{-0.0625(5)}\approx0.731616"

ii.

The memoryless property says that


"P(X>r+t|X>r)=P(X>t)"

"P(X>5+5|X>5)=P(X>5)"

"=e^{-0.0625(5)}\approx0.731616"